# Phase diagram and Coarsening dynamics in dry apolar Active Nematics with Reciprocal local alignment

### v1

@@ Abstract
Using the Lebwohl-Lasher interaction for reciprocal local alignment, we present a comprehensive phase diagram for a dry, apolar, active nematic system using its stochastic dynamics. The nematic-isotropic transition in this system is first-order and fluctuation-dominated. Our phase diagram identifies three distinct regions based on activity and orientational noise relative to alignment strength: a homogeneous isotropic phase, a nematic phase with giant density fluctuations, and a coexistence region. Using mean-field analysis and hydrodynamic theory, we demonstrate that reciprocal interactions lead to a density fluctuation-induced first-order transition and derive a phase boundary consistent with numerical results. Quenching from the isotropic to nematic phase reveals coarsening dynamics with distinctive scaling behaviors for density and nematic fields, exhibiting exponents between 2 and 3, typical of non-conserved and conserved dynamics.
@@ 1 Introduction
Active nematics represent a fascinating class of nonequilibrium systems characterized by the collective motion and organization of self-propelled entities [1, 2, 3, 4]. Inspired by a myriad of biological systems, ranging from the dynamic organization of the cytoskeleton within cells to the emergent behaviors observed in tissues and even granular materials, active nematics have emerged as a pivotal area of study at the interface of physics, biology, and materials science [4, 5, 3, 6, 7]. At the heart of active systems lies the concept of self-propulsion, where individual constituents continually convert internal energy into directed motion. This intrinsic activity, prevalent in biological systems such as motile cells and bacterial colonies, fuels a rich array of dynamic phenomena, including the spontaneous formation of orientational order, collective motion, and the emergence of complex spatiotemporal patterns [4, 5, 6]. Moreover, active nematics find analogs in nonliving systems, including vibrated granular matter and artificial microswimmers suspended in fluids. The last few decades have seen tremendous progress in understanding their collective properties and phase transitions [1, 2, 3, 4, 8, 9, 10].
@@
In dry active matter, self-propelled particles (SPP) with local alignment display order-disorder transition coupled with density fluctuations. Consideration of ferromagnetic alignment of the active heading directions of individual propulsion led to flocking and persistent motion of the flocks. This was demonstrated in the celebrated Vicsek model and its variants and described by the Toner-Tu theory of coupled orientation and density fields [11, 12, 13, 14, 4]. Numerical studies of polar rods with volume exclusion showed nematic alignment and clustering [15], and a related simplified model led to system-spanning clusters [16]. A Boltzmann-Landau-Ginzburg kinetic theory approach, based on a Vicsek-like model of active nematics with non-reciprocal torques, produced hydrodynamic equations consistent with previous findings [2, 3, 4, 8, 9, 17]. In certain systems, alignment results from physical processes such as actual collisions between active elements [18, 15, 19]. Other instances involve effective interactions mediated by non-equilibrium mechanisms like complex biochemical signaling or visual and cognitive processes [20, 21, 22]. The former alignments are characterized by reciprocal interactions, while non-reciprocal rules often describe the latter [23, 24, 25]. In equilibrium systems, microscopic details are typically considered irrelevant to emergent macroscopic properties if the models share symmetries and conservation laws. However, recent studies on aligning active particles with short-range interactions, be it vectorial or nematic active matter, revealed that microscopic implementations in the form of reciprocity and additivity can influence macroscopic behaviors [20], e.g., leading to first-order or critical order-disorder transition in active apolar nematics in the presence or absence of reciprocity [26].
@@
SPPs can lose their overall polarity when there are rapid reversals in the self-propulsion direction. In active nematics, particles spontaneously align along an axis $\hat{\bf n}$ with a $\hat{\bf n}\to-\hat{\bf n}$ symmetry [27, 2, 17, 28, 29, 10, 30]. Examples of such systems include colliding elongated objects [15, 16, 31], migrating cells [32, 33], cytoskeletal filaments [34], certain direction-reversing bacteria [35, 36, 37, 38, 39], and vibrated granular rods [18, 19]. Even at high activity levels, the apolar nature of the particles results in zero macroscopic velocity. Intriguingly, the phase behavior of these apolar SPPs still shows a clear dependence on activity [28, 8, 9, 17, 10].
@@
In this paper, we consider a dry, apolar, active nematics model using a reciprocal additive alignment interaction between local neighbors. An earlier on-lattice implementation using the Lebwohl-Lasher potential of alignment [40] found a fluctuation-dominated phase separation associated with the ordering transition [28]. Our earlier off-lattice calculation of SPPs at a fixed activity demonstrated a clear first-order nematic-isotropic (NI) transition with increasing orientational noise [26]. In this paper, we present a comprehensive phase diagram as a function of changing activity and relative orientational noise with respect to the alignment strength. It shows three different parameter regions characterized by (i) a homogeneous isotropic phase, (ii) a nematically aligned phase with significant density fluctuations, and (iii) a coexistence of nematic and isotropic in the presence of the largest density fluctuations. The order-disorder transition is fluctuation-dominated and first-order. Using a mean-field hydrodynamic argument, we obtain an analytic prediction for the phase boundary that shows reasonable agreement with our numerical results. Moreover, we study the coarsening dynamics associated with phase ordering after a quench from isotropic to nematic phase. The stochastic simulations show distinctive dynamical scaling for density and nematic fields, with clearly distinguishable scaling exponents, unlike the earlier predictions from mean-field hydrodynamics [41].
@@
The paper is structured as follows: Section 2 details the model of dry active nematic with reciprocal alignment. Section 3 presents the comprehensive phase diagram. Section 4 describes the mean field hydrodynamic analysis, explaining the phase boundaries and the nature of the nematic-isotropic transition. We then explore the coarsening of nematic order and density field in Section 5. Finally, in Section-6, we conclude by summarizing our main results.
@@ 2 Model
Here, $\mu$ represents mobility, $D_{r}$ is the rotational diffusion constant, and $dB_{i}$ is a Gaussian process with zero mean and correlation $\langle dB_{i}dB_{j}\rangle=\delta_{ij}dt$. These equations depict persistent motion for a free particle, where $D_{r}$ determines the persistence time $\tau_{p}=D_{r}^{-1}$, setting the unit of time. The model illustrates apolar particles aligning nematically in the presence of interaction ($J>0$). It is important to note that the torque experienced by a particle pair $i,j$ is equal and opposite to each other.
Employing the Euler-Maruyama scheme, we perform direct numerical simulations by integrating Eq.(1) and (3) using $dt=7\times 10^{-4}\,D_{r}^{-1}$. Care is taken so that $v_{0}dt<r_{c}$ in order to incorporate the influence of interaction properly. We utilize scaled angular diffusion $D_{\theta}:={2D_{r}}/{\mu J}$ and activity $Pe:=v_{0}/D_{r}r_{c}$ as control parameters to construct the phase diagram, delineating associated phases at a fixed dimensionless number density $\rho r_{c}^{2}={Nr_{c}^{2}}/{L^{2}}=0.4$.
From numerical simulations, we assess the degree of nematic order utilizing the scalar order parameter:
@@
where the average is taken over all $i$ particles. The value of $S$ is bounded on both sides, $0\leq S\leq 1$. To compute the steady-state average $S$ across the entire system, we average over all $N$ particles and further average over steady-state configurations, represented by red line points in Fig.1(b). Additionally, we determine the standard deviation of the global order parameter $S$ and identify its maximum as the transition point $D_{\theta}^{\ast}$ for a given $Pe$; these are indicated by blue line points in Fig.1(b). For a localized assessment of $S$, we perform averaging over particles within a local volume. This is used to determine $S$ over coarse-grained volumes of linear dimension $10r_{c}$ to calculate the probability distributions $P(S)$ in Fig.1(c).
@@ 3 Phase diagram
The three insets show the typical distributions of the scalar nematic order parameter corresponding to nematic (N), coexistence (C), and isotropic (D) regions. The black solid line indicates a $D_{\theta}\sim Pe^{2}$ boundary between the isotropic and nematic phases. An analytic description of the NI phase boundary $D_{\theta}\sim Pe^{2}$ is provided in the following section. The broken solid line indicates a breakdown of this simple estimate. As expected, the nematic phase is characterized by a giant number fluctuation [26, 28]. Deep inside the nematic phase, e.g., at $Pe=428.57$ and $D_{\theta}=0.28$ the system shows $\Delta n^{2}=\langle n^{2}\rangle-\langle n\rangle^{2}\sim n^{a}$ with $a\approx 1.8$. The steady-state pair correlation function $C(r)=1-b(r/L)^{\alpha}$ with $\alpha\approx 0.26$ [26, 42] signifies a clear violation of Porod’s law in the coexistence region. This captures a huge roughness of interfaces evidenced by large $(1-\alpha)$ characterizing a fluctuation-dominated phase separation.
The NI transition is captured using the dependence of the scalar order parameter $S$ and its standard deviation $\sigma_{s}=[\langle S^{2}\rangle-\langle S\rangle^{2}]^{1/2}$ on $D_{\theta}$, at a fixed $Pe$; see Fig.1(b). The local probability distribution $P(S)$ in Fig.1(c) shows an unimodal distribution with the maximum at small $S$ for large $D_{\theta}$, transforms into a bimodal distribution at intermediate $D_{\theta}$ values. This clearly indicates a phase coexistence, characterizing a first-order transition. At an even smaller $D_{\theta}$, the distribution becomes unimodal again, this time displaying a single maximum at large $S$, indicating a nematic phase.

### v2

@@ Abstract
Using the Lebwohl-Lasher interaction for reciprocal local alignment, we present a comprehensive phase diagram for a dry, apolar, active nematic system using its stochastic dynamics. The nematic-isotropic transition in this system is first-order and fluctuation-dominated. Our phase diagram identifies three distinct regions based on activity and orientational noise relative to alignment strength: a homogeneous isotropic phase, a nematic phase with giant density fluctuations, and a coexistence region. Using mean-field analysis and hydrodynamic theory, we demonstrate that reciprocal interactions lead to a density fluctuation-induced first-order transition and derive a phase boundary consistent with numerical results. Quenching from the isotropic to nematic phase reveals coarsening dynamics with nematic ordering preceding particle clustering and similar scaling behaviors for density and nematic fields, exhibiting dynamic exponents between 2 and 3.
@@ 1 Introduction
Active nematics represent a fascinating class of nonequilibrium systems characterized by the collective motion and organization of self-propelled entities [1, 2, 3, 4, 5, 6]. Inspired by a myriad of biological systems, ranging from the dynamic organization of the cytoskeleton within cells to the emergent behaviors observed in tissues and even granular materials, active nematics have emerged as a pivotal area of study at the interface of physics, biology, and materials science [4, 7, 3, 8, 9]. At the heart of active systems lies the concept of self-propulsion, where individual constituents continually convert internal energy into directed motion. This intrinsic activity, prevalent in biological systems such as motile cells and bacterial colonies, fuels a rich array of dynamic phenomena, including the spontaneous formation of orientational order, collective motion, and the emergence of complex spatiotemporal patterns [4, 7, 8]. Moreover, active nematics find analogs in nonliving systems, including vibrated granular matter and artificial microswimmers suspended in fluids. The last few decades have seen tremendous progress in understanding their collective properties and phase behaviors [1, 2, 3, 4, 10, 11, 12, 13].
@@
In dry active matter, self-propelled particles (SPP) with local alignment display order-disorder transition coupled with density fluctuations. Consideration of ferromagnetic alignment of the active heading directions of individual propulsion led to flocking and persistent motion of the flocks. This was demonstrated in the celebrated Vicsek model and its variants and described by the Toner-Tu theory of coupled orientation and density fields [14, 15, 16, 17, 4]. Numerical studies of polar rods with volume exclusion showed nematic alignment and clustering [18], and a related simplified model led to system-spanning clusters [19]. A Boltzmann-Landau-Ginzburg kinetic theory approach, based on a Vicsek-like model of active nematics, produced hydrodynamic equations consistent with previous findings [2, 3, 4, 10, 11, 20]. In certain systems, alignment results from physical processes such as actual collisions between active elements [21, 18, 22]. Other instances involve effective interactions mediated by non-equilibrium mechanisms like complex biochemical signaling or visual and cognitive processes [23, 24, 25]. The former alignments are characterized by reciprocal interactions, while non-reciprocal rules often describe the latter [26, 27, 28]. In equilibrium systems, microscopic details are typically considered irrelevant to emergent macroscopic properties if the models share symmetries and conservation laws. However, recent studies on aligning active particles with short-range interactions, be it vectorial or nematic active matter, revealed that microscopic implementations in the form of reciprocity and additivity can influence macroscopic behaviors [23], e.g., leading to first-order or critical order-disorder transition in active apolar nematics in the presence or absence of reciprocity [29].
@@
SPPs can lose their overall polarity when there are rapid reversals in the self-propulsion direction. In active nematics, particles spontaneously align along an axis $\hat{\bf n}$ with a $\hat{\bf n}\to-\hat{\bf n}$ symmetry [1, 2, 20, 30, 31, 12, 32]. Examples of such systems include colliding elongated objects [18, 19, 33], migrating cells [34, 35], cytoskeletal filaments [36], certain direction-reversing bacteria [37, 38, 39, 40, 41], and vibrated granular rods [21, 22]. Even at high activity levels, the apolar nature of the particles results in zero macroscopic velocity. Intriguingly, the collective properties of these apolar SPPs still show a clear dependence on activity, as orientational fluctuations drive particle current [2, 30, 10, 11, 20, 12].
@@
In this paper, we consider a dry, apolar, active nematics model using a reciprocal additive alignment interaction between local neighbors. An earlier on-lattice implementation using the Lebwohl-Lasher potential of alignment [42] found a fluctuation-dominated phase separation associated with the ordering transition [30]. Our earlier off-lattice calculation of SPPs at a fixed activity demonstrated a clear first-order nematic-isotropic (NI) transition with increasing orientational noise [29].
In this paper, we present a comprehensive phase diagram as a function of changing activity and relative orientational noise with respect to the alignment strength. It shows three different parameter regimes characterized by (i) a homogeneous isotropic phase, (ii) a nematically aligned phase with significant density fluctuations, and (iii) a coexistence of nematic and isotropic in the presence of the largest density fluctuations. The order-disorder transition remains fluctuation-dominated and first-order across all parameter regimes. Using a mean-field hydrodynamic argument, we obtain an analytic prediction for the phase boundary that shows reasonable agreement with our numerical results. Our primary contribution is this direct calculation of the detailed phase diagram.
Moreover, we study the phase ordering dynamics after a deep quench from isotropic to nematic phase. The stochastic simulations show dynamical scaling for density and nematic fields, with similar scaling exponents with values bounded between non-conserved and conserved dynamics consistent with earlier hydrodynamic calculations [43]. We observe a delay in particle clustering dynamics relative to nematic ordering, with the delay time controlled by the depth of the quench. Our second main contribution is the direct calculation of phase ordering using stochastic simulations.
@@
The rest of the paper is structured as follows: Section 2 details the model of dry active nematic with reciprocal alignment. Section 3 presents the comprehensive phase diagram. Section 4 describes the mean field hydrodynamic analysis, explaining the phase boundaries and the nature of the nematic-isotropic transition. We then explore the coarsening of nematic order and density field in Section 5. Finally, in Section 6, we conclude by summarizing our main results and presenting an outlook.
@@ 2 Model
Here, $\mu$ represents mobility, $D_{r}$ is the rotational diffusion constant, and $dB_{i}$ is a Gaussian process with zero mean and correlation $\langle dB_{i}dB_{j}\rangle=\delta_{ij}dt$. These equations depict persistent motion for a free particle, where $D_{r}$ determines the persistence time $\tau_{p}=D_{r}^{-1}$, setting the unit of time. The model illustrates apolar particles aligning nematically with strength $J>0$. It is important to note that the torque experienced by a particle pair $i,j$ is equal and opposite to each other.
Employing the Euler-Maruyama scheme, we perform direct numerical simulations by integrating Eq.(1) and (3) using $dt=7\times 10^{-4}\,D_{r}^{-1}$. Care is taken so that $v_{0}dt<r_{c}$ in order to incorporate the influence of interaction properly. We utilize scaled angular diffusion $D_{\theta}:={2D_{r}}/{\mu J}$ and activity $Pe:=v_{0}/D_{r}r_{c}$ as control parameters to construct the phase diagram, delineating associated phases at a fixed dimensionless number density $\rho={N}/{L^{2}}=\rho_{0}$ with $\rho_{0}r_{c}^{2}=0.4$.
The nematic order parameter is $\langle Q_{\alpha\beta}\rangle$ with $Q^{(i)}_{\alpha\beta}=(u_{i\alpha}u_{i\beta}-\frac{1}{d}\delta_{\alpha\beta})$ for $i$-th particle with $\alpha,\beta$ denoting the components. In 2D,
${\bf Q}^{(i)}=\frac{1}{2}\begin{pmatrix}\cos 2\theta_{i}&\sin 2\theta_{i}\\ \sin 2\theta_{i}&-\cos 2\theta_{i}\end{pmatrix}.$
From numerical simulations, we assess the degree of nematic order by utilizing the positive eigenvalue of $\langle Q_{\alpha\beta}\rangle$, the scalar order parameter:
@@
The value of $S$ is bounded on both sides, $0\leq S\leq 1$. To compute the steady-state average $S$ across the entire system, we average over all $N$ particles and further average over steady-state configurations, represented by red line points in Fig.1(b). Additionally, we determine the standard deviation of the global order parameter $S$ and identify its maximum as the transition point $D_{\theta}^{\ast}$ for a given $Pe$; these are indicated by blue line points in Fig.1(b). For a localized assessment of $S$, we perform averaging over particles within a local volume. This is used to determine $S$ over coarse-grained volumes of linear dimension $10r_{c}$ to calculate the probability distributions $P(S)$ in Fig.1(c).
@@ 3 Phase diagram
The three insets show the typical distributions of the scalar nematic order parameter corresponding to nematic (N), coexistence (C), and isotropic (D) regions. An analytic description of the NI phase boundary $D_{\theta}\sim Pe^{2}$ is provided in the following section. Shown by a black solid line in Fig.1(a), it qualitatively captures the actual phase boundary up to $Pe\approx 300$. The broken solid line indicates a breakdown of this simple estimate. Beyond this point, the phase boundary becomes approximately independent of $Pe$. As expected, the active nematic phase is characterized by a giant number fluctuation [29, 30, 2]. For example, deep inside the nematic phase, at $Pe=714.28$ and $D_{\theta}=0.254$, the system shows $\langle\Delta n\rangle^{2}=\langle n^{2}\rangle-\langle n\rangle^{2}\sim\langle n\rangle^{a}$ with $a\approx 1.52$; see Appendix-A. The steady-state pair correlation function $C(r)=1-b(r/L)^{\alpha}$ with $\alpha\approx 0.26$ [29] signifies a clear violation of Porod’s law in the coexistence region. This captures a huge roughness of interfaces evidenced by large $(1-\alpha)$ characterizing a fluctuation-dominated phase separation [44].
The NI transition is captured using the dependence of the scalar order parameter $S$ and its standard deviation $\sigma_{s}=[\langle S^{2}\rangle-\langle S\rangle^{2}]^{1/2}$ on $D_{\theta}$, at a fixed $Pe$; see Fig.1(b). The cross-correlation between fluctuations in nematic order and density peaks near the transition, as detailed in Appendix-B. However, this correlation diminishes significantly within the nematic phase and approximately vanishes after transitioning to isotropic. The local probability distribution $P(S)$ in Fig.1(c) shows an unimodal distribution with the maximum at small $S$ for large $D_{\theta}$, transforms into a bimodal distribution at intermediate $D_{\theta}$ values. This clearly indicates a phase coexistence, characterizing a first-order transition. At an even smaller $D_{\theta}$, the distribution becomes unimodal again, this time displaying a single maximum at large $S$, indicating a nematic phase.
@@ 4 Mean field and hydrodynamic analysis
Here, we consider the coupled evolution of the nematic order parameter density $\Pi_{ij}({\bf r},t)=Q_{ij}({\bf r},t)\rho({\bf r},t)$ and particle density $\rho({\bf r},t)$ [10, 11, 20, 4]. The local field of nematic order
${\bf Q}({\bf r},t)=$ $\frac{S({\bf r},t)}{2}\begin{pmatrix}\cos 2\theta({\bf r},t)&\sin 2\theta({\bf r},t)\\ \sin 2\theta({\bf r},t)&-\cos 2\theta({\bf r},t)\end{pmatrix}$ (5)
is determined by the local scalar order $S({\bf r},t)$ and orientation $\theta({\bf r},t)$. Using active current $J^{(a)}_{i}=-Pe\,\partial_{j}\Pi_{ji}$ [4], the particle density field evolves as,
where $c_{i}$ denotes components of random current taken to be statistically isotropic and delta correlated, and $D$ denotes the diffusion constant. The evolution of nematic order follows,
where $q_{ij}$ is an uncorrelated Gaussian tensorial noise, which is traceless and symmetric, denoting thermal or active fluctuations [4]. Here, the coefficients $\alpha_{1}=[1-{D_{\theta}}/{D_{\theta}^{(c)}}]$ and $\alpha_{2}=({1}/{4})[{D_{\theta}^{(c)}}/{D_{\theta}}]^{2}$ with the equilibrium mean-field critical point $D_{\theta}^{(c)}=A\rho$ [29]; see Appendix-C.
In the homogeneous limit of constant $\rho$, Eq.(7) gives $\frac{dQ_{ij}}{dt}=[\alpha_{1}-\alpha_{2}{\rm Tr}(Q^{2})]Q_{ij}$ leading to a continuous transition at $\alpha_{1}=0$. Describing the first-order transition requires incorporating the density fluctuations [45, 46], for which we use a renormalized mean field theory [47, 32]. We consider a small activity-induced density fluctuation $\delta\rho$ such that $\rho=\rho_{0}+\delta\rho$ to express $\alpha_{1,2}(\rho)=\alpha_{1,2}^{0}+{\alpha^{\prime}}_{1,2}^{0}\delta\rho$ where $\alpha_{1,2}^{0}=\alpha_{1,2}(\rho_{0})$ and ${\alpha^{\prime}}_{1,2}^{0}=(\partial\alpha_{1,2}/\partial\rho)_{\rho_{0}}$. The requirement of a zero current steady state in density evolution suggests $\delta\rho\approx PeS/D$ [29]. Using this in the mean-field limit of Eq.(7), we obtain the evolution for scalar order
$\partial_{t}S=[u_{2}+u_{3}S-u_{4}S^{2}]S,$ (8)
where, $u_{2}=\alpha_{1}^{0}$, $u_{3}={\alpha^{\prime}}_{1}^{0}Pe/D$ and $u_{4}=\alpha_{2}^{0}$. This equation can be expressed as a free energy minimizing kinematics $\partial_{t}S=-\partial{\cal F}/\partial S$ with
${\cal F}=-\frac{u_{2}}{2}S^{2}-\frac{u_{3}}{3}S^{3}+\frac{u_{4}}{4}S^{4},$ (9)
keeping up to $S^{4}$ order term.
Figure 3: The nematic and density autocorrelation $C_{\bf Q}(r)$ ($a$) and $C_{\rho}(r)$ ($b$) during quench from $D_{\theta}=2.8$ to $D_{\theta}=0.28$ at $Pe=428.57$. Insets show the correlations plotted against $r/L$. The main plots show data collapse by scaling separation $r$ by dynamic correlation lengths $\xi(t)$ ($a$) and $\zeta(t)$ ($b$). $(c)$ The correlation lengths $\xi(t)$ ($\square$), $\zeta(t)$ ($\circ$) in the scaling regime exhibit similar power-law growths $\sim t^{1/z}$ with exponent $z\approx 2.5$, indicated by the solid green line. The growth exponent $32<z<$ 3 is shown using the dashed blue line of $t^{1/2}$ scaling and the dash-dotted black line of $t^{1/3}$ scaling. Cells of linear dimension $5\,r_{c}$ are used to calculate the coarse-grained correlations from numerical simulations.
The cubic term in $S$ in the effective free energy density ${\cal F}$ results in a first-order NI transition. At the transition, the scalar order parameter jumps from $S=0$ to
$S_{c}=\frac{2}{3}\frac{{\alpha^{\prime}}_{1}^{0}}{\alpha_{2}^{0}}\frac{Pe}{D}.$ (10)
The transition point changes with $Pe$ to give the transition line
increasing quadratically with $Pe$. A plot of $D_{\theta}^{\ast}\sim Pe^{2}$ using the black solid line in Fig.1(a) approximately captures the phase boundary up to $D_{\theta}\approx 1.4$. At higher $D_{\theta}$, even the activity-induced increased density fluctuations fail to sustain the nematic order, rendering the phase boundary independent of $Pe$. In fact, as can be observed from the heat map in Fig.1(a), the density fluctuation gets suppressed at these $D_{\theta}$ values. We note that the first-order NI transition is purely active; it vanishes in the limit of $Pe\to 0$ with the vanishing of $S_{c}$.
Before closing this section, certain comments are in order. Within the Lebowhl-Lasher model used here, the first-order isotropic-nematic transition precludes the possible continuous transition as $D_{\theta}^{\ast}>D_{\theta}^{(c)}$. Moreover, the observed first-order transition relies crucially on the density dependence of $D_{\theta}^{(c)}$, which resulted from the reciprocal and additive nature of interaction implemented in the Lebowhl-Lasher model. If $D_{\theta}^{\ast}$ is independent of density, as in certain non-reciprocal models, the NI transition can become continuous [29].
5 Phase ordering kinematics
Having established the phenomenology of the NI transition, we now aim to understand the coarsening kinematics of the nematic order. Towards this end, we present our numerical results of the growth kinetics following a deep quench at a fixed $Pe=428.57$ from a homogeneous isotropic initial state at $D_{\theta}=2.8$ to $D_{\theta}=0.28$ corresponding to a steady-state nematic. For better statistics, we use a larger system with $N=64000$. Fig.2 shows a typical series of snapshots depicting this evolution. As time progresses, domains of higher order parameters and higher density form, with their typical size increasing over time. The deep quench leads to local instability. In the beginning, the homogeneous isotropic state shows instability through the formation of nematically ordered filaments crisscrossing each other; see Fig.2(b). The small filaments merge and coarsen with time, as shown in Fig.2(c) and (d). At a late time, one finds dynamic system-spanning bands of nematically ordered regions; see Fig.2(e) and (f). In the steady state, the local undulations in nematic bands make them unstable toward breaking, and one finds repeated formation, breaking, and reorientations of such bands with time [12, 11].
To quantify these observations, we track the spatial nematic order parameter autocorrelation $C_{\bf Q}(r,t)$ and the density autocorrelation $C_{\rho}(r,t)$ as the system relaxes towards its final ordered state. To obtain a direct comparison against earlier field theory calculations, we divide the simulation volume into cells of length $\ell_{c}$. We employ the coarse-grained order parameter ${\bf Q}({\bf r},t)=\frac{1}{n(t)}\sum_{i=1}^{n(t)}{\bf Q}^{(i)}$ by performing instantaneous averaging over local cells at positions ${\bf r}$ containing $n(t)$ number of particles. Similarly, in the same cells, we use a coarse-grained local density $\rho({\bf r},t)=n(t)/\ell_{c}^{2}$. We use $\ell_{c}=5r_{c}$ in the numerical calculations. The correlation functions are defined as follows:
$C_{\bf Q}(r,t)=\frac{Tr\langle{\bf Q}(r,t){\bf Q}(0,t)\rangle}{Tr\langle{\bf Q}^{2}(0,t)\rangle},$ (12)
and
where $\delta\rho(r)=\rho(r)-\rho_{0}$ and $r$ denotes the separation between local coarse-grained volumes. For further details, see Appendix-D.
The time-dependent correlation functions plotted in Fig.3 are calculated averaging over forty independent trajectories. The simulations show exponentially decaying correlations up to $tD_{r}\approx 500$. Beyond that, the finite-size effect kicks in; the deep quench renders quasi-long ranged nematic order with power-law decay of correlations and system-spanning bands. At shorter times, the correlation length $\ell$ can be obtained by fitting the correlation functions with $\exp(-r/\ell)$ form. Rescaling the separations by correlation lengths $\ell=\xi(t)$ and $\ell=\zeta(t)$ for nematic and density correlations, we obtain nice data collapse for both the correlation functions, shown in the main figures of Fig.3($a$) and ($b$).
In the intermediate time scaling regime, both the correlation lengths, $\zeta(t)$ and $\xi(t)$, increase algebraically. Note that the building up of orientational correlation precedes that of local density; see the delay in growth of $\zeta(t)$ compared to $\xi(t)$ in Fig.3($c$). This delay is caused by the time required for nematic instability to influence the density fluctuation. As can be seen from Eq.(6), the typical delay time is controlled by the quenched value of $D_{\theta}$ itself.
In the scaling regime, the correlation lengths grow following power laws $\xi(t)\sim{t}^{1/z_{S}}$ and $\zeta(t)\sim t^{1/z_{\rho}}$, with dynamic exponents $z_{S},z_{\rho}\approx 2.5$; see Fig.3($c$). It is interesting to note that earlier equilibrium simulations of a soft-spin version of the 2D Lebowhl-Lasher model led to an approximate phase ordering exponent $z_{S}\approx 2.5$ [48] differing from $\xi\propto(t/\ln t)^{1/2}$ growth law expected for equilibrium nematic [49]. In our simulations, the value of $z_{\rho}$ is expected to be related to $z_{S}$ due to the density-order coupling. Indeed, the bounds of non-conserved and conserved dynamics [50] to these exponents, $32<z_{S},z_{\rho}<$ 3, are consistent with earlier non-stochastic hydrodynamic estimates [43]. A precise determination of the growth laws and growth exponents requires performing a more careful and extensive finite-size scaling procedure.
6 Outlook
We presented a comprehensive phase diagram as a function of activity $Pe$ and relative orientational noise $D_{\theta}$ for the nematic-isotropic transition in dry active nematics with reciprocal alignment interaction. The additive nature of the interaction ensures a first-order transition via coupling of nematic order to density fluctuations, confirmed through a hydrodynamic theory and the corresponding mean-field arguments.
It is intriguing that despite zero macroscopic velocity, the phase behavior of these apolar SPPs shows intricate dependence on $Pe$. Our theoretical analysis predicts a phase boundary $D_{\theta}\sim Pe^{2}$, aligning well with simulation results up to $D_{\theta}\approx 1.4$. Beyond this, local density increases are not enough to sustain nematic order, resulting in a $Pe$-independent phase boundary. The scalar order parameter shows a discontinuous jump proportional to $Pe$ at the transition, disappearing in the passive particle limit of vanishing $Pe$. We identified a broad parameter regime of phase coexistence. Deep within the nematic phase, the nematic bands are inherently unstable, exhibiting giant fluctuations and turbulent dynamics due to their continuous formation and breakup.
Moreover, we investigated phase ordering kinetics by quenching from the isotropic to the nematic phase. This revealed coarsening patterns with correlation lengths in nematic order and density fluctuations growing as $t^{1/z}$ with exponents, $z_{S},z_{\rho}\approx 2.5$ remaining within the bounds of non-conserved and conserved dynamics, $32<z_{S},\,z_{\rho}<$ 3, a range consistent with earlier hydrodynamic predictions in Ref. [43]. The quench in scaled orientational coupling showed a lag in particle clustering relative to the growth of nematic correlation.
In conclusion, we obtained a comprehensive phase diagram for reciprocal alignment in active nematics, supported by mean-field predictions for the observed transition and phase boundaries. A theoretical description beyond the mean field explaining the full phase diagram will require further work. The numerical studies of coarsening dynamics revealed dynamic exponents with values in between the exponents for non-conserved and conserved dynamics. A precise numerical determination of growth laws requires careful system size scaling studies. The extent to which the observed phenomenology holds for non-reciprocal interactions warrants further investigation.
Author Contributions
DC designed and supervised the research. AS performed all the numerical calculations and data analysis. DC and AS wrote the manuscript.
Figure 4: The number fluctuation $\langle\Delta n^{2}\rangle$ as a function of mean number $\langle n\rangle$ at $Pe=714.28$ and $D_{\theta}=0.254$, calculated over square coarse graining areas of linear dimension $\ell_{c}/r_{c}=4,5,8,10,16,20,25,40,50$. A fitting, the solid line, shows $\langle\Delta n^{2}\rangle\sim\langle n\rangle^{a}$, with $a\approx 1.52$.
Acknowledgments
The numerical simulations were performed using SAMKHYA, the High-Performance Computing Facility provided by the Institute of Physics, Bhubaneswar. D.C. thanks Sriram Ramaswamy and Abhishek Chaudhuri for useful discussions and acknowledges the Department of Atomic Energy, Government of India (1603/2/2020/IoP/R&D-II/150288) and the International Centre for Theoretical Sciences (ICTS-TIFR), Bangalore, for an Associateship.
Appendix A Giant number fluctuations
Deep inside the nematic phase, at $Pe=714.28$ and $D_{\theta}=0.254$, we analyze the particle number fluctuations $\langle\Delta n^{2}\rangle=\langle n^{2}\rangle-\langle n\rangle^{2}$ as a function of $\langle n\rangle$; see Fig.4. This shows giant number fluctuations, characterized by $\langle\Delta n^{2}\rangle\sim\langle n\rangle^{a}$, with $a\approx 1.52>1$ but less than $a=2$ predicted in [2] and found by numerical simulations in [30, 31]. See Fig.1($g$) for a typical configuration corresponding to these parameter values.
Figure 5: Cross-correlaion coefficient $C_{\rho S}$ with the scaled orientation noise $D_{\theta}$ at a fixed $Pe=428.57$.
Appendix B Density-nematic order cross correlation
Here, we compute the cross-correlation $C_{\rho S}=\langle\delta\rho(r)\delta S(r)\rangle/\sigma_{S}\sigma_{\rho}$, where $\delta\rho(r)=\rho(r)-\langle\rho\rangle$ and $\delta S(r)=S(r)-\langle S\rangle$ with $\langle\rho\rangle$ and $\langle S\rangle$ denoting the mean density and mean scalar order, respectively(see Fig.5). Further, $\sigma_{S}=\langle\delta S^{2}(r)\rangle^{1/2}$ and $\sigma_{\rho}=\langle\delta\rho^{2}(r)\rangle^{1/2}$ denote the standard deviations of nematic and density fields. The calculations are performed over coarse-grained cells of linear dimension $10\,r_{c}$. In the disordered phase, $D_{\theta}>1.5$, the cross-correlation $C_{\rho S}$ is slightly negative, indicating an anti-correlation between the fields. This happens because an increase in randomly oriented elements within a coarse-grained region can lead to better cancellation in averaging and reduce the overall nematic order. Near the transition point, $D_{\theta}\approx 1.4$, $C_{\rho S}$ reaches its positive maximum because higher particle density in bands enhances order, while the low-density background remains disordered (see Fig.1($f$) ). Deeper in the nematic phase, $D_{\theta}\lesssim 0.28$, $C_{\rho S}$ decreases as regions of both high and low densities exhibit similar nematic order (see Fig. 1($g$) ).
Appendix C Equilibrium mean-field theory for nematic-isotropic transition